# Practical and Accurate Local Edge Differentially Private Graph Algorithms

### v2

@@ 1. Introduction
Consider the application of $k$-core decomposition to COVID transmission data (Qin et al., 2020; S Monteiro et al., 2021; Serafino et al., 2022; Guo et al., 2022; Chang and Ferrara, 2022) and other disease transmission networks (Ciaperoni et al., 2020) such as HIV (Fujimoto et al., 2022; Holme and Litvak, 2017). The core numbers are generated and published, sometimes even with location data (Serafino et al., 2020). Revealing the precise core numbers of every individual can lead to privacy breaches. Consider a scenario where exactly $c$ individuals have a core number of $c-1$. This implies they form a clique of $c$ vertices, all connected. In disease transmission graphs, this directly exposes a cluster of sensitive disease transmissions! Therefore, it’s essential to release privacy-preserving core numbers.
@@
Table 1. Additive error bounds compared to previous work. Here $n:$ number of nodes, $D_{max}:$ maximum degree, $\varepsilon:$ privacy parameter, $d:$ degeneracy, and $C_{4}$ and $\overrightarrow{C}_{4}$ is the number of $4$-cycles and oriented $4$-cycles, respectively.
@@ 2. Preliminaries
[…]
@@ 2.1. Privacy Tools
Let $\mathcal{M}$ be an $\varepsilon$-LEDP algorithm and $h$ be an arbitrary (randomized) mapping from $Range(\mathcal{M})$ to an arbitrary set. The algorithm $h\circ\mathcal{M}$ is $\varepsilon$-LEDP.444$\circ$ is notation for applying $h$ on the outputs of $\mathcal{M}$.
@@ 2.2. Problem Definitions
Figure 2. Example $k$-core decomposition and triangles in a $4$-degenerate graph. Nodes are assigned core numbers based on the highest value core they belong to; e.g., a node in the $1$-core but not in the $2$-core is given the core number of $1$. Larger valued cores are contained within all smaller valued cores; e.g., the $3$-core is contained in the $1$ and $2$-core. Red edges show the triangles, i.e., $3$-cycles in the graph. The degeneracy of this graph is $4$.
@@ 6       Set $\mathcal{C}\leftarrow$ new $\text{Coordinator}\left(LDS,channel\right)$
10             maxDegreeThresholds$\left[w\right]=\textnormal{{DegreeThresholdWorker}}(w,\varepsilon_{1},L,b)$
@@ 14             Set $\mathcal{F}(r)\leftarrow\left\lfloor\frac{r}{L}\right\rfloor$
[…]
@@
18            $\mathcal{C}.\text{wait}()$triangleright coordinator waits for workers to finish
@@
21                   if nextLevels$\left[i\right]=1$ then
@@ 2 Function DegreeThresholdWorker($w,\varepsilon,L,b$)
3       Set maxThreshold0absent0$\leftarrow$ 0← 0
4       for node ${}v:=$ localGraph do
@@ 5             Sample $X\sim\mathsf{Geom}\left(\frac{\varepsilon}{2}\right)$
6             $\widetilde{d}_{v}\leftarrow d_{v}+X$triangleright noised degree
@@
8             $v$.threshold $\leftarrow\left\lceil{\log_{2}(\widetilde{d}_{v})}\right\rceil\cdot L$triangleright thresholding
@@
1 Input: worker id $w$; round number $r$; privacy parameter $\varepsilon\in(0,1]$; constant $\psi$; group index $\mathcal{F}(r)$; pointer to the coordinator $LDS$.
2 Function LevelMovingWorker($w,r,\varepsilon,\psi,\mathcal{F}(r),LDS$)
@@ 3       Set nextLevels $\leftarrow[0,\ldots,0]$
4       for node ${}v:=$ localGraph do
5             if v.threshold $=r$ then
6                   $v$.permZero = 00
@@
10             if vLevel $=r$ and v.permZero $\neq 0$ then
@@ 14                   Set extra bias $B\leftarrow\frac{6e^{s}}{\left(e^{2s}-1\right)^{3}}$
17                         nextLevels$\left[v\right]=1$
@@ 3.2. Theoretical Analysis
Let $M$ be the number of workers and $n$ the graph size. Each worker processes $S$ nodes, where $S=\lfloor n/M\rfloor$ for $M-1$ workers, and the last worker handles $n-(M-1)\lfloor n/M\rfloor$. The coordinator maintains the level data structure (LDS) and a communication channel, both requiring $O(n)$ space, resulting in a total memory usage of $O(n)$. Each worker processes $O(S)$ nodes, requiring $O(Sn)$ space for the graph and an additional $O(S)$ space for auxiliary structures, leading to a total of $O(Sn)$. In terms of communication, workers send one bit per node per round, incurring a per-worker cost of $O(S)$ and an overall round cost of $O(n)$. The coordinator receives and distributes the updated LDS, adding another $O(n)$ cost. Thus, the total communication overhead for the algorithm is $O\left(n\log(n)\log\left(D_{\max}\right)\right)$.
@@ 8             $\textnormal{{RRWorker}}\left(w,n,\frac{\varepsilon}{4}\right)$
10      $\mathcal{C}.\text{wait}()$triangleright coordinator waits for workers to finish
@@ 12       $\triangleright$ Round 2: Max Out-degree
[…]
@@ 18       $\triangleright$ Round 3: Count Triangles
[…]
@@ 2 Function RRWorker($w,n,\varepsilon$)
3       Set neighborsRR [][$\leftarrow$[][]← [ ] [ ]
4       for node ${}v:=$ localGraph do
@@ 5             For all $ngh_{v}\leftarrow\left\{j:j\in\left[n\right]\wedge j>v\right\}$
6             neighborsRR$\left[v\right]=\text{RandomizedResponse}_{\varepsilon}(ngh_{v})$
@@ 2 Function MaxOutDegreeWorker($w,\varepsilon,Z$)
4       for node, adjacency list ${}v,ngh:=$ localGraph do
@@ 2 Function CountTrianglesWorker($w,\varepsilon,Z,X,\widetilde{d}_{\max}$)
4       for node, adjacency list ${}v,ngh:=$ localGraph do
@@ 4.1. Algorithm Description
[…]
@@ 4.2. Theoretical Analysis
Let $M$ be the number of workers and $n$ the graph size. Each worker processes $S$ nodes, where $S=\lfloor n/M\rfloor$ for $M-1$ workers, and the last worker handles $n-(M-1)\lfloor n/M\rfloor$ nodes. The coordinator manages three communication channels and publishes the noisy edges for the entire graph. The cRR structure, which aggregates noisy edges, requires $O(n^{2})$ space, while cTCount and cMaxOut, which collect triangle counts and maximum noisy out-degree, require $O(M)$ space each. Storing published noisy edges further adds $O(n^{2})$ space, resulting in a total coordinator memory requirement of $O(n^{2}+M)$. Each worker processes $O(S)$ nodes, requiring $O(Sn)$ space for the graph. The neighborsRR structure for storing noisy edges demands $O(Sn)$ space, while computing the maximum noisy out-degree requires $O(S)$. The final triangle count computation takes $O(S\cdot\widetilde{d}_{\max})$ space, where $\widetilde{d}_{\max}$ is the maximum noisy out-degree, leading to an overall worker memory requirement of $O(Sn)$. The algorithm runs three communication rounds beyond those for low out-degree ordering. Workers first send noisy edges, incurring $O(Sn)$ communication cost, followed by sending the maximum noisy out-degree and triangle counts, each requiring $O(M)$ communication. Thus, the total communication overhead for the algorithm is $O(n^{2}+M)$.
@@
Figure 5. Oriented cycle of length $4$; two non-adjacent black nodes have edges oriented toward the remaining red nodes.
@@
We first prove that, in expectation, $\mathbb{E}\left[\frac{X_{j,k}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}\right]=\mathbbm{1}_{j,k}$ where $\mathbbm{1}_{j,k}$ is the indicator variable for whether edge $\{j,k\}$ exists in the original private graph. Since randomized response flips the bit indicating the existence of an edge between each pair of vertices with probability $\frac{1}{e^{\varepsilon/4}-1}$, we can simplify the expression to be:
@@
We must account for the symmetric geometric noise. Since the expectation of the symmetric geometric noise is 00 and we add together all of the values for each of the drawn noises, the expected total noise is 00 by linearity of expectations. Each triangle has a unique node that queries it since every triangle has a unique orientation of edges where two edges are outgoing from a vertex in the triangle. Hence, the expected sum of all queried triangles is $T$, conditioned on $U$, by linearity of expectations.
@@
Now, it remains to compute $\mathrm{Var}\left[\sum_{v\in[n]}\sum_{j,k\in Out(v)}X_{j,k}\right]$. The covariance is 00 if two queried pairs do not query the same $X_{j,k}$. The covariance is non-zero only in the case of queries which overlap in $X_{i,j}$. This occurs only in the case of an oriented $4$-cycle where $X_{i,j}$ is shared between two queries. Let $P_{2}$ be the set of all pairs of such queries that share $X_{j,k}$. In this case, the covariance is upper bounded by $\mathbb{E}[X_{j,k}^{2}]$. Hence, we can simplify our expression to be
@@ 5. Experimental Evaluation
Parameters We use $\varepsilon=1.0$, bias term $=8$, approximation factor $(2+\eta)=5.625$ (matching non-private $k$-core experiments (Liu et al., 2022b)), and privacy split fraction $f=0.8$ (allocating $0.8\cdot\varepsilon$ to thresholding and $0.2\cdot\varepsilon$ to level moving step). We also conduct an ablation study on key parameters $\varepsilon$ and $f$, and our theoretical proofs show the approximation falls within a $(2+\eta)$-multiplicative factor.
@@ 5.2. Triangle Counting
On small or dense graphs like wiki or email-eu-core, the TCountRR—which incurs a dominant error term growing proportional to $\frac{n^{3/2}}{\varepsilon^{2}}$—performs similarly to ours because $n$ is small (Eden et al., 2023a). However, in larger graphs, $n^{3/2}$ explodes, whereas our error depends instead on $\sqrt{n}\,d$ (with $d\ll n$). TCountRR must also consider all $O(n^{4})$ $4$-cycles, but our method only needs $\widetilde{O}(n^{2}d^{2})$ $4$-cycles. Consequently, on larger datasets such as enron and brightkite, we reduce the relative error (Fig. 10(b)) by roughly 17x and 45.5x, respectively. Compared to using 60 worker cores, we achieve up to a $1.41$x speedup.

### v3

@@ 1. Introduction
Consider the application of $k$-core decomposition to COVID transmission data (Qin et al., 2020; S Monteiro et al., 2021; Serafino et al., 2022; Guo et al., 2022; Chang and Ferrara, 2022) and other disease transmission networks (Ciaperoni et al., 2020) such as HIV (Fujimoto et al., 2022; Holme and Litvak, 2017). The core numbers are generated and published, sometimes even with location data (Serafino et al., 2020). Revealing the precise core numbers of every individual can lead to privacy breaches. Consider a scenario where exactly $c$ individuals have a core number of c1c-1. This implies they form a clique of $c$ vertices, all connected. In disease transmission graphs, this directly exposes a cluster of sensitive disease transmissions! Therefore, it’s essential to release privacy-preserving core numbers.
@@
Table 1. Additive error bounds compared to previous work. Here $n:$ number of nodes, $D_{max}:$ maximum degree, $\varepsilon:$ privacy parameter, $d:$ degeneracy, and $C_{4}$ and $\overrightarrow{C}_{4}$ is the number of $4$-cycles and oriented 44-cycles, respectively.
@@ 2. Preliminaries
Algorithm $\mathcal{A}(G)$, that takes as input a graph $G$ and outputs some value in $Range(\mathcal{A})$, [^3] Range()Range($\cdot$) denotes the set of all possible outputs of a function. is $\varepsilon$-edge differentially private ($\varepsilon$-edge DP) if for all $S\subseteq Range(\mathcal{A})$ and all edge-neighboring graphs $G$ and $G^{\prime}$,
@@ 2.1. Privacy Tools
[…]
@@ 2.2. Problem Definitions
Figure 2. Example $k$-core decomposition and triangles in a 44-degenerate graph. Nodes are assigned core numbers based on the highest value core they belong to; e.g., a node in the $1$-core but not in the $2$-core is given the core number of $1$. Larger valued cores are contained within all smaller valued cores; e.g., the $3$-core is contained in the $1$ and 22-core. Red edges show the triangles, i.e., 33-cycles in the graph. The degeneracy of this graph is $4$.
@@ 8  Set maxDegreeThresholds $\leftarrow\left[\right]$
10     maxDegreeThresholds[$\left[w\right]=\textnormal{{DegreeThresholdWorker}}(w,\varepsilon_{1},L,b)$
@@ 14     Set $\mathcal{F}(r)\leftarrow\left\lfloor\frac{r}{L}\right\rfloor$
16        $\textnormal{{LevelMovingWorker}}\left(w,r,\varepsilon_{2},\psi,\mathcal{F}(r),\mathcal{C}.LDS\right)$
@@
18    $\mathcal{C}.\text{wait}()$ $\triangleright$ coordinator waits for workers to finish
@@
21        if nextLevels[$\left[i\right]=1$ then
@@ 2 Function DegreeThresholdWorker($w,\varepsilon,L,b$)
[…]
4  for node v:=~v:= localGraph do
@@ 5     Sample $X\sim\mathsf{Geom}\left(\frac{\varepsilon}{2}\right)$
6     $\widetilde{d}_{v}\leftarrow d_{v}+X$ $\triangleright$ noised degree
@@
8     $v$.threshold $\leftarrow\left\lceil{\log_{2}(\widetilde{d}_{v})}\right\rceil\cdot L$ $\triangleright$ thresholding
@@
1 Input: worker id $w$; round number $r$; privacy parameter $\varepsilon\in(0,1]$; constant $\psi$; group index $\mathcal{F}(r)$; pointer to the coordinator LDSLDS.
[…]
@@ 3  Set nextLevels $\leftarrow[0,\ldots,0]$
4  for node v:=~v:= localGraph do
5     if v.threshold =r=r then
6        $v$.permZero = 0
@@
10     if vLevel =r=r and v.permZero $\neq 0$ then
@@ 14        Set extra bias $B\leftarrow\frac{6e^{s}}{\left(e^{2s}-1\right)^{3}}$
17           nextLevels[$\left[v\right]=1$
@@ 3.2. Theoretical Analysis
Let $M$ be the number of workers and $n$ the graph size. Each worker processes $S$ nodes, where $S=\lfloor n/M\rfloor$ for M1M-1 workers, and the last worker handles $n-(M-1)\lfloor n/M\rfloor$. The coordinator maintains the level data structure (LDS) and a communication channel, both requiring $O(n)$ space, resulting in a total memory usage of $O(n)$. Each worker processes $O(S)$ nodes, requiring $O(Sn)$ space for the graph and an additional $O(S)$ space for auxiliary structures, leading to a total of $O(Sn)$. In terms of communication, workers send one bit per node per round, incurring a per-worker cost of $O(S)$ and an overall round cost of $O(n)$. The coordinator receives and distributes the updated LDS, adding another $O(n)$ cost. Thus, the total communication overhead for the algorithm is $O\left(n\log(n)\log\left(D_{\max}\right)\right)$.
@@ 8     $\textnormal{{RRWorker}}\left(w,n,\frac{\varepsilon}{4}\right)$
10 $\mathcal{C}.\text{wait}()$ $\triangleright$ coordinator waits for workers to finish
@@ 12  $\triangleright$ Round 2: Max Out-degree
14     $\textnormal{{MaxOutDegreeWorker}}\left(w,\frac{\varepsilon}{4},Z\right)$
@@ 18  $\triangleright$ Round 3: Count Triangles
20     $\textnormal{{CountTrianglesWorker}}\left(w,\frac{\varepsilon}{4},Z,X,\widetilde{d}_{\max}\right)$
@@ 2 Function RRWorker($w,n,\varepsilon$)
3  Set neighborsRR $\leftarrow[][]$
4  for node v:=~v:= localGraph do
@@ 5     For all $ngh_{v}\leftarrow\left\{j:j\in\left[n\right]\wedge j>v\right\}$
6     neighborsRR[$\left[v\right]=\text{RandomizedResponse}_{\varepsilon}(ngh_{v})$
@@ 2 Function MaxOutDegreeWorker($w,\varepsilon,Z$)
4  for node, adjacency list v,ngh:=~v,ngh:= localGraph do
@@ 2 Function CountTrianglesWorker($w,\varepsilon,Z,X,\widetilde{d}_{\max}$)
4  for node, adjacency list v,ngh:=~v,ngh:= localGraph do
@@ 4.1. Algorithm Description
Worker (Randomized Response) As specified in Algorithm 6, workers maintain a neighborsRR data structure to store noisy edges. For each node $v$, noisy edges are computed via Randomized Response (RR) with parameter $\varepsilon$, processing only the upper triangular part of the adjacency matrix (Algorithm 6), as the graph is undirected. Specifically, for a node $v$, all indices greater than $v$ are processed using $\text{RandomizedResponse}_{\varepsilon}(ngh_{v})$, which flips the existence of each edge $(v,ngh_{v})$ with probability $\frac{1}{e^{\varepsilon}+1}$ (Algorithm 6). Once computed, workers send noisy edges to the coordinator (Algorithm 6).
@@ 4.2. Theoretical Analysis
Let $M$ be the number of workers and $n$ the graph size. Each worker processes $S$ nodes, where $S=\lfloor n/M\rfloor$ for M1M-1 workers, and the last worker handles $n-(M-1)\lfloor n/M\rfloor$ nodes. The coordinator manages three communication channels and publishes the noisy edges for the entire graph. The cRR structure, which aggregates noisy edges, requires $O(n^{2})$ space, while cTCount and cMaxOut, which collect triangle counts and maximum noisy out-degree, require $O(M)$ space each. Storing published noisy edges further adds $O(n^{2})$ space, resulting in a total coordinator memory requirement of $O(n^{2}+M)$. Each worker processes $O(S)$ nodes, requiring $O(Sn)$ space for the graph. The neighborsRR structure for storing noisy edges demands $O(Sn)$ space, while computing the maximum noisy out-degree requires $O(S)$. The final triangle count computation takes $O(S\cdot\widetilde{d}_{\max})$ space, where $\widetilde{d}_{\max}$ is the maximum noisy out-degree, leading to an overall worker memory requirement of $O(Sn)$. The algorithm runs three communication rounds beyond those for low out-degree ordering. Workers first send noisy edges, incurring $O(Sn)$ communication cost, followed by sending the maximum noisy out-degree and triangle counts, each requiring $O(M)$ communication. Thus, the total communication overhead for the algorithm is $O(n^{2}+M)$.
@@
Figure 5. Oriented cycle of length 44; two non-adjacent black nodes have edges oriented toward the remaining red nodes.
@@
We first prove that, in expectation, $\mathbb{E}\left[\frac{X_{j,k}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}\right]=\mathbbm{1}_{j,k}$ where $\mathbbm{1}_{j,k}$ is the indicator variable for whether edge $\{j,k\}$ exists in the original private graph. Since randomized response flips the bit indicating the existence of an edge between each pair of vertices with probability $\frac{1}{e^{\varepsilon/4}+1}$, we can simplify the expression to be:
@@
We must account for the symmetric geometric noise. Since the expectation of the symmetric geometric noise is 0 and we add together all of the values for each of the drawn noises, the expected total noise is 0 by linearity of expectations. Each triangle has a unique node that queries it since every triangle has a unique orientation of edges where two edges are outgoing from a vertex in the triangle. Hence, the expected sum of all queried triangles is $T$, conditioned on $U$, by linearity of expectations.
@@
Now, it remains to compute $\mathrm{Var}\left[\sum_{v\in[n]}\sum_{j,k\in Out(v)}X_{j,k}\right]$. The covariance is 0 if two queried pairs do not query the same $X_{j,k}$. The covariance is non-zero only in the case of queries which overlap in $X_{i,j}$. This occurs only in the case of an oriented $4$-cycle where $X_{i,j}$ is shared between two queries. Let $P_{2}$ be the set of all pairs of such queries that share $X_{j,k}$. In this case, the covariance is upper bounded by $\mathbb{E}[X_{j,k}^{2}]$. Hence, we can simplify our expression to be
@@ 5. Experimental Evaluation
Parameters We use $\varepsilon=1.0$, bias term =8=8, approximation factor $(2+\eta)=5.625$ (matching non-private $k$-core experiments (Liu et al., 2022b)), and privacy split fraction $f=0.8$ (allocating $0.8\cdot\varepsilon$ to thresholding and $0.2\cdot\varepsilon$ to level moving step). We also conduct an ablation study on key parameters $\varepsilon$ and $f$, and our theoretical proofs show the approximation falls within a $(2+\eta)$-multiplicative factor.
@@ 5.2. Triangle Counting
On small or dense graphs like wiki or email-eu-core, the TCountRR—which incurs a dominant error term growing proportional to $\frac{n^{3/2}}{\varepsilon^{2}}$—performs similarly to ours because $n$ is small (Eden et al., 2023a). However, in larger graphs, $n^{3/2}$ explodes, whereas our error depends instead on $\sqrt{n}\,d$ (with $d\ll n$). TCountRR must also consider all $O(n^{4})$ 44-cycles, but our method only needs $\widetilde{O}(n^{2}d^{2})$ 44-cycles. Consequently, on larger datasets such as enron and brightkite, we reduce the relative error (Fig. 10(b)) by roughly 17x and 45.5x, respectively. Compared to using 60 worker cores, we achieve up to a 1.411.41x speedup.